# APCodec: A Neural Audio Codec with Parallel Amplitude and Phase Spectrum Encoding and Decoding

### v1

@@ IV-E2 Validation on Diverse Audio Datasets
Since the VCTK is a speech dataset, to assess the performance of the proposed APCodec on different types of audio datasets, we incorporated two additional datasets. These two additional datasets included: Opencpop [47], a publicly available high quality Mandarin singing corpus of approximately 5.2 hours designed for singing voice synthesis; FSD50K [48], an open dataset of approximately 391 hours for human-labeled sound events. For the Opencpop dataset, we utilized the officially pre-trimmed data, selecting 3,367 utterances as the training set and the remaining 389 utterances as the test set. For the FSD50K dataset, 40,966 utterance and 10,231 utterances were respectively chosen as the training set and test set. The sampling rates for the Opencpop and FSD50K datasets are both 44.1 kHz. We upsampled them to 48 kHz for experiments.
@@
For the sake of fairness and simplicity, we compared the performance between non-streamable HiFi-Codec and APCodec, as well as the performance between streamable AudioDec and APCodec-S, respectively. These models were further finetuned for 200k steps each on the Opencpop and FSD50K datasets, based on the well-trained models using the VCTK dataset. We separately calculated the objective metrics for the Opencpop and FSD50K datasets on their respective test sets, and the results are shown in Table IV. Due to the fact that the STOI is typically used solely for assessing speech intelligibility, we exclusively employed the LSD, ViSQOL, and $\text{AWPD}_{\text{IP}}$ metrics in this experiment. It can be observed that, whether on the Opencpop or the FSD50K dataset, the performance for all metrics of the APCodec was significantly better than that of the HiFi-Codec. Specifically, on the sound event dataset (i.e., FSD50K), the performance of the HiFi-Codec noticeably declined compared to its performances on datasets like VCTK and Opencpop, which involve human vocalizations. However, our proposed APCodec still managed to maintain good performance. In comparison between the AudioDec and APCodec-S on the Opencpop dataset, the gap between the LSD of the AudioDec and the LSD of the APCodec-S significantly narrowed, compared to the results on the VCTK dataset. The AudioDec even led by 0.04 on the ViSQOL score. However, on the FSD50K dataset, all metrics of the AudioDec were inferior to those of the APCodec-S. The above experimental results fully demonstrated that our proposed APCodec exhibited strong generalization and adaptability on other types of datasets, especially on non-human vocalization datasets, compared to other mainstream neural codecs. Thus, the APCodec is more suitable for various audio signal processing tasks which will also be a focus of our future work.
V Conclusion
@@
In this paper, we proposed a novel neural audio codec called APCodec. The APCodec leveraged the advantages of parametric codecs, regarding the audio amplitude and phase spectra as parametric characteristics rather than the raw waveforms for parallel encoding and parallel decoding. Thus, it could obtain latent codes at low sampling rate using very minimal downsampling operations. To ensure the fidelity of the decoded audio similar to waveform codecs, spectral-level loss, quantization loss, and GAN-based loss were employed to train the APCodec model. We also constructed a low-latency streamable APCodec by combining feed-forward layers and causal convolution layers with knowledge distillation training strategies. Experimental results confirm that our proposed APCodec exhibited advantages at high waveform sampling rates and low bitrates, demonstrating high-quality decoded audio, low compression rate, fast generation speed, low model complexity, and low latency. It surpassed the performance of the baseline SoundStream, Encodec, HiFi-Codec and AudioDec. Further analysis experiments also confirmed the effectiveness of the structure and loss proposed in APCodec, as well as its versatility and generalizability across diverse audio datasets.

### v2

@@ III-E2 Validation on Diverse Audio Datasets
Since the VCTK is a small-scale speech dataset, to assess the performance of the proposed APCodec on different sizes and types of audio datasets, we incorporated three additional datasets. These three additional datasets included: Common Voice [49], a large-scale massively-multilingual transcribed speech corpus of approximately 919 hours; Opencpop [50], a publicly available high quality Mandarin singing corpus of approximately 5.2 hours designed for singing voice synthesis; FSD50K [51], an open dataset of approximately 84 hours for human-labeled sound events. For the Common Voice dataset [^7] https://commonvoice.mozilla.org/en/datasets., we used the “Common Voice Corpus 17.0” data on the download website and selected speech utterances with a sampling rate of 48 kHz. 568,822 and 6,026 utterances were respectively chosen as the training set and test set. For the Opencpop dataset [^8] https://wenet.org.cn/opencpop/., we utilized the officially pre-trimmed data, selecting 3,367 utterances as the training set and the remaining 389 utterances as the test set. For the FSD50K dataset [^9] https://zenodo.org/records/4060432., 40,966 and 4,436 utterances were respectively chosen as the training set and test set. The sampling rates for the Opencpop and FSD50K datasets are both 44.1 kHz. We upsampled them to 48 kHz for experiments.
@@
For the sake of fairness and simplicity, we compared the performance between non-streamable APCodec and DAC, as well as the performance between streamable APCodec-S and AudioDec, respectively. These models were further finetuned for 200k steps each on the Common Voice, Opencpop and FSD50K datasets, based on the well-trained models using the VCTK dataset. We separately calculated the objective metrics for these three datasets on their respective test sets, and the results are shown in Table IV. Due to the fact that the STOI is typically used solely for assessing speech intelligibility, we exclusively employed the LSD, ViSQOL, and $\text{AWPD}_{\text{IP}}$ metrics in this experiment. It can be observed that, whether on the Common Voice, Opencpop or the FSD50K dataset, the performance for all metrics of the APCodec was significantly better than that of the DAC. This confirms that our proposed APCodec still outperformed the DAC on larger speech datasets or other types of audio datasets. When comparing the APCodec-S and AudioDec on the Common Voice and Opencpop datasets, despite the APCodec-S being superior in terms of amplitude and phase quality, its overall objective quality, according to ViSQOL results, was slightly inferior to that of AudioDec. However, on the FSD50K dataset, all metrics of the AudioDec were inferior to those of the APCodec-S. The above experimental results fully demonstrated that our proposed APCodec exhibited strong generalization and adaptability on other types of datasets, especially on non-human vocalization datasets, compared to other mainstream neural codecs. Thus, the APCodec is more suitable for various audio signal processing tasks which will also be a focus of our future work.
IV Conclusion
@@
In this paper, we proposed a novel neural audio codec called APCodec. The APCodec leveraged the advantages of parametric codecs, regarding the audio amplitude and phase spectra as parametric characteristics rather than the raw waveforms for parallel encoding and parallel decoding. Thus, it could obtain latent codes at low frame rate using very minimal downsampling operations. To ensure the fidelity of the decoded audio similar to waveform codecs, spectral-level loss, quantization loss, and GAN-based loss were employed to train the APCodec model. We also constructed a low-latency streamable APCodec by combining feed-forward layers and causal deconvolutional layers with knowledge distillation training strategies. Experimental results confirm that our proposed APCodec exhibited advantages at high waveform sampling rates and low bitrates, demonstrating high-quality decoded audio, high compression rate, fast generation speed, low model complexity, and low latency. It surpassed the performance of the baseline Encodec, AudioDec and DAC. Further analysis experiments also confirmed the effectiveness of the structure and loss proposed in APCodec, as well as its versatility and generalizability across diverse audio datasets.